One year ago, Jamie purchased a house valued at $480,000. To avoid the insurance (PMI), she decided to pay down 21%. She got a 30-year fully amortized fixed rate mortgage at 4.75% interest rate and her closing costs and discount points were about $5,600 (around 1.5% of loan amount). 2. What is the APR on Jamie’s mortgage?
Answer #1

Answer:

Loan amount = Value of house * (1 - down payment %) = 480000 * (1 - 21%) = $379,200

Given:

Number of months = 30 * 12 = 360

Monthly interest rate = 4.75%/12

Monthly payment = PMT(rate, nper, pv, fv, type) = PMT (4.75%/12, 360, -379200, 0, 0) = $1978.0867

Since Jamie paid closing costs and discount points amounting to = $5,600

Actual amount she received = 379200 - 5600 = $373,600

Now we calculate monthly interest rate using RATE function with PV as = $373,600

= RATE (nper, pmt, pv, fv, type) = RATE (360, 1978.0867, -373600, 0, 0) = 0.4066%

Hence:

APR = 0.4066% * 12 = 4.88%

APR on Jamie’s mortgage = 4.88%